1. Solve by substitution or elimination method:
3x – 2y = 26
-7x + 3y = -49
2. Solve by substitution or elimination method:
4x – 5y = 14
-12x + 15y = -42
3. Solve by substitution or elimination method:
-2x + 6y = 19
10x – 30y = -15

Second set of problems:
1. Plot the graph of the equations 2x - 3y = 5 and 4x - 6y = 21 and interpret the result.
2. Plot the graph of the equations 4x - 6y = 12 and x - 5y = 10 and interpret the result.
3. Plot the graph of the equations 8x - 4y = 12 and -4x + 2y = -6 and interpret the result.

Answer by RAY100(1637)

I),,(1) 3x-2y=26,,mult by 3,,,9x-6y= 78
(2)-7x +3y =-49,,,mult by 2,,-14x+6y=-98
add(1) +(2),,,,,-5x=-20,,,,,x=4
subst into (1),,,3(4) -2y=26,,,,-2y=14,,,,y=-7
.
II),,,(1)4x-5y=14,,,mult by (-3),,,-12x +15y =-42
(2) -12x +15y =-42
as (1) and (2) are coincident,,,,x= all real numbers
.
III) (1) -2x -6y = 19,,,mult by (-5),,,10x-30y=-95
(2) 10x-30y=-15
as (1) and (2) are parallel(slopes same),,,there is no solution
.
(iv) (1)2x-3y=5,,,y= (2/3)x -5/3
(2) 4x-6y=21,,,,,y=(2/3)x -7/2
lines are parallel(slopes same), y intercepts are (-5/3) and (-7/2)
.
(v) (1) 4x-6y=12,,y=(2/3)x-2
(2) x-5y=10,,,y=(1/5)x -2
y intercepts are same,,therefore lines intercept there(0,-2),,
both slopes are positive, but different.
.
(vi) (1) 8x -4y =12,,,,y=2x-3
(2) -4x +2y =-6,,,y=2x-3
lines are coincident,,,y intercept is -3,,(0,-3),,slope is (+2)
.